# Robust $\mathcal{H}_{\infty}$ control and worst-case search in constrained parametric space

### v2

@@ V-B Closed loop and control problem
The benchmark focuses on the attitude control of the satellite described above to follow a reference $r$. The control input $u$ of the plant $G(s,\delta)$ is the torque applied to the satellite (e.g. by reaction wheels, thrusters…) and the output vector contains the 3 angle measurements polluted by some noise $n$. The closed loop is represented in Figure 2, where $d(s)$ represents avionics: a 2-nd order Pade approximation of a 0.125s delay, and a 2-nd order low-pass filter with a cut-off frequency of 10 rad/s representing actuator dynamics. We note $\mathcal{K}(s)$ the set of controllers $K(s)$ of the desired structure: one 4-th order controller per spacecraft axis, each one taking the measured angle and rate as inputs and generating a control torque, resulting in 24 decision parameters per controller (72 in total). The closed-loop plant has 50 states.
@@ V-C Worst-case search
In this section, we present the worst-case search for the spectral abscissa and the $\mathcal{H}_{\infty}$ norm, as presented in Section III. We selected three controllers $K_{1}$, $K_{2}$, $K_{3}$ obtained at early iterations of the robust controller tuning presented in Section V-D and analyze the robustness of the closed-loop system with three Test cases. The tests are ran on a 12th Gen Intel(R) Core(TM) i7-12700H 2.30 GHz, with 14 cores and 16GB of RAM. Computation times are provided for reference, but they may vary depending on hardware, software configuration, implementation details, optimization parameters, etc.
V-C1 Test case 1: spectral abscissa (multiple local optima)
Test case 1 illustrates how the proposed approaches perform in presence of several local optima, in different regions of the parametric space, that are not particularly rare. We search for $\max_{\delta\in\mathcal{D}}\;\alpha(A(K_{1},\delta))$.
@@ V-C1 Test case 1: spectral abscissa (multiple local optima)
The parameters of the Particle Swarm Optimization (PSO) were chosen as follows to favor global exploration over local exploitation. We set a range of 0.6 to 1.1 for the adaptive inertias (’InertiaRange’ option in MATLAB). Iterations terminate when performance is increased by less than 10% (”FunctionTolerance’) across 20 last iterations (’MaxStallIterations’). For a given computational budget, because PSO is sensitive to initialization [Qian2020], running several independent instances with different initializations can yield better results than increasing the number of particles or tightening the termination tolerance. Therefore, we performed 100 independent runs with the parameters given above, each with 500 particles (’SwarmSize’) – cf. also [piotrowski2020] for a discussion on population size in PSO. All other parameters were set to MATLAB’s default settings. Bearing in mind that our parameter choices were specifically tuned for multiple fast independent runs rather than to maximize the reliability of a single run, the results should be interpreted to determine how many such runs are required to reliably identify a worst case. We see that, with a few runs, we are able to consistently identify the worst-case configuration that was most often missed by the local SQP alone.
@@
V-C2 Test case 2: spectral abscissa (rare worst case)
Test case 2 illustrates how the proposed approaches perform when the destabilization is rare. Indeed, only $\sim$ 0.018% of the admissible parametric space (as approximated by the MC experiments) yields unstability of $A(K_{2},\delta)$. We search for $\max_{\delta\in\mathcal{D}}\;\alpha(A(K_{2},\delta))$.
(a) SQP, with 100 random initializations
(b) Poles map (zoomed) for a SQP run
(c) PSO and PSO+SQP, with random 100 initializations
(d) MC and MC+SQP, repeated 100 times
Figure 4: Worst-case spectral abscissa for Test case 2
Most SQP runs remained stuck on the local optimum $\alpha(A(\delta^{*}_{\text{SQP}}))=-a(\delta^{*}_{\text{SQP}})=-2\mathrm{e}^{-7}$ that occupies most of the parametric space (Figure 4(a)). On the other hand, PSO+SQP (Figure 4(c)) more consistently finds the destabilizing configurations. PSO runs required more function evaluations than Test case 1: 1525500 across the 100 runs, for a runtime ranging from 6s to 57s with an average of 11s. MC (again executed with the same total number of objective function evaluations as PSO for fair comparison) is also able to detect unstability (Figure 4(d)), but it is significantly outperformed by PSO (compare the blue crosses of Figures 4(c) and 4(d)). MC+SQP performs decently but is slightly less reliable than PSO+SQP.
V-C3 Test case 3: $\mathcal{H}_{\infty}$ norm
[…]
The results are shown in Figure 5. The SQP runs took between 38s and 1172s (average: 591s). The PSO individual runtimes ranged from 72 to 126s (average: 85s), and MC around 95s on average. Here, PSO tends to get stuck in the local optimum 1.15, whereas MC+SQP more frequently converges to the suggested global optimum. In both cases, the subsequent SQP improves the result of PSO or MC alone.
(a) SQP, with 100 random initializations
(b) Sigma plot: 100 random samples and a few KKT points
(c) PSO and PSO+SQP, with random 100 initializations
(d) MC and MC+SQP, repeated 100 times
Figure 5: Worst-case (normalized) $\mathcal{H}_{\infty}$ norm for Test case 3
@@ V-D Robust control
In this section, we solve Problem (26) using Algorithm 2.
@@ VI Conclusion
To perform the worst-case search, we combined Particle Swarm Optimization (or Monte-Carlo sampling, if stochastic guarantees are desired) for global exploration and Sequential Quadratic Programming for local exploitation. Numerical experiments on a flexible spacecraft benchmark show that this method significantly outperforms Monte-Carlo sampling alone, and is able to reliably detect even rare worst-case configurations. Then, by iteratively constructing a set of active configurations, multi-objectives robust controller optimization was performed in reasonable runtime.
Finally, we emphasize that the proposed worst-case search, performing global exploration and being compatible with a large number of uncertain parameters and states, has the potential to complement existing robust control techniques, including in the unconstrained case.
@@ Acknowledgments
The authors acknowledge the use of an AI tool (ChatGPT: GPT-4) to assist in the reformulation of certain sentences.
Appendix A Appendix: Mathematical proofs
In this Appendix, we prove the propositions of Section II. For readability, they are reminded before introducing the proofs.
[…]
Proof.
The proof is provided in the body of the paper. ∎
a) Descent direction. Let $(p,u)$ be a KKT pair of $Q(x,H,g_{x})$. If $p||\neq 0$ and $u||_{\infty}\leq r$, then $p$ is a descent direction for $\theta_{r}(x)$, i.e. $\theta^{\prime}_{r}(x,p)<0$ where $\theta^{\prime}_{r}(x,d)$ denotes the the directional derivative at $x$ in direction d.
Proof.
a) The proof of a) is an adaptation of [Han1997, Theorem 3.1], originally written for a differentiable $f$, to the upper-$C^{1}$ case. Let us define:
$I=\{i:c_{i}(x)>0\},\;\bar{I}=\{i:c_{i}(x)=0\},^{;}\hat{I}=\{i:c_{i}(x)<0\}$
As in [Han1997] we have:
$\theta^{\prime}_{r}(x,p)=f^{\prime}(x,p)+r\sum_{i\in I}\nabla c_{i}(x)^{T}p+r\sum_{i\in\bar{I}}(\nabla c_{i}(x)^{T}p)_{+}\;.$
Since $f$ is upper-$C^{1}$, we have $f^{\prime}(x,p)=\min_{g\in\partial f(x)}\langle g,p\rangle\leq g_{x}^{T}p$ and thus
$\theta^{\prime}_{r}(x,p)\leq g_{x}^{T}p+r\sum_{i\in I}\nabla c_{i}(x)^{T}p+r\sum_{i\in\bar{I}}(\nabla c_{i}(x)^{T}p)_{+}\;.$
In the original proof of [Han1997, Theorem 3.1], the expression above is an equality with $\nabla f(x)$ instead of $g_{x}$. From there, the proof of [Han1997, Theorem 3.1] is identical, as follows. Since $(p,u$) is a KKT pair of $Q(x,H,g_{x})$, we have $c_{i}(x)+\nabla c_{i}(x)^{T}p\leq 0$ which yields:
$\sum_{i\in\bar{I}}(\nabla c_{i}(x)^{T}p)_{+}=0$
and since $u_{i}(c_{i}(x)+\nabla c_{i}(x)^{T}p)=0$, we deduce:
$\theta^{\prime}_{r}(x,p)\leq g_{x}^{T}p+\sum_{i=1}^{n}u_{i}\nabla c_{i}(x)^{T}p+\sum_{i=1}^{n}u_{i}c_{i}(x)+r\sum_{i\in I}\nabla c_{i}(x)^{T}p\;.$
Since $(p,u$) is a KKT pair of $Q(x,H,g_{x})$ we also have:
$g_{x}+\sum_{i=1}^{n}u_{i}\nabla c_{i}(x)+\frac{1}{2}(H+H^{T})p=0$
and noticing that $\sum_{i\in\bar{I}\cup\hat{I}}^{n}u_{i}c_{i}(x)\leq 0$ yields:
$\theta^{\prime}_{r}(x,p)$ $\leq-\frac{1}{2}(H+H^{T})p+\sum_{i\in I}^{n}(c_{i}(x)+r\nabla c_{i}(x)^{T}p)$
$\leq-\frac{1}{2}(H+H^{T})p+\sum_{i\in I}^{n}(u_{i}-r)c_{i}(x)$
$<0$
since $H$ is positive definite and $|u|_{\infty}<r$.
b) is proved in the body of the paper.
To prove Proposition 3, we first introduce the following lemmas.
Lemma 1 (Supermonotonicity of the subdifferential of upper-$C^{1}$ functions)
Let $f$ be upper-$C^{1}$ at $x$. For every $\epsilon>0$, there exists $\delta>0$ such that for all $x_{0},x^{\prime}_{0}\in\mathcal{B}(x,\delta)$:
[…]
where $g_{x_{0}}\in\partial f(x_{0})$ and $g_{x^{\prime}_{0}}\in\partial f(x^{\prime}_{0})$.
Proof.
This comes from [Spingarn81, Theorem 3.9] (submonotonicity of the subdifferential of lower-$C^{1}$ functions) and is also noted for example in [Apkarian2016bundle, Lemma 6]. ∎
Lemma 2 (Perturbation of a quadratic problem) Let us consider the two quadratic problems:
$x_{0}=\arg\min\left\{\;Q(x)=\frac{1}{2}x^{T}Kx-x^{T}k\;\;\text{, subject to }\;Gx\leq g\right\}$
$x^{\prime}_{0}=\arg\min\;\left\{Q^{\prime}(x)=\frac{1}{2}x^{T}K^{\prime}x-x^{T}k^{\prime}\;\;\text{, subject to }\;G^{\prime}x\leq g^{\prime}\right\}$
[…]
$x_{0}^{\prime}-x_{0}||\leq c\epsilon$
whenever $\epsilon<\epsilon_{0}$.
Proof.
[…]
$x_{0}=\arg\min\left\{Q(x)=\frac{1}{2}x^{T}Kx-x^{T}k\right\}\;,\quad x^{\prime}_{0}=\arg\min\left\{Q^{\prime}(x)\frac{1}{2}x^{T}K^{\prime}x-x^{T}k^{\prime}\right\}$
we have:
[…]
The proof for this statement is almost identical to the one provided in [Daniel1973, Theorem 2.1], as follows. For any $x\in\mathbb{R}^{k}$, we have:
$(x-x_{0})^{T}\nabla Q(x_{0})\geq 0$
$(x-x^{\prime}_{0})^{T}\nabla Q^{\prime}(x^{\prime}_{0})\geq 0$
with the gradients $\nabla Q(x)=Kx-k$ and $\nabla Q(x)=K^{\prime}x-k^{\prime}$. Substituting $x=x^{\prime}_{0}$ and $x=x_{0}$ respectively in these two equations and adding the results yields
$(x^{\prime}_{0}-x_{0})^{T}(\nabla Q(x_{0})-\nabla Q^{\prime}(x^{\prime}_{0}))\geq 0$
which is equivalent to
$(x^{\prime}_{0}-x_{0})^{T}(\nabla Q^{\prime}(x^{\prime}_{0})-\nabla Q^{\prime}(x_{0}))\leq(x^{\prime}_{0}-x_{0})^{T}(\nabla Q(x_{0})-\nabla Q^{\prime}(x_{0}))$
For the left side:
$(x^{\prime}_{0}-x_{0})^{T}(\nabla Q^{\prime}(x^{\prime}_{0})-\nabla Q^{\prime}(x_{0}))$ $=(x^{\prime}_{0}-x_{0})^{T}K^{\prime}(x^{\prime}_{0}-x_{0})$
[…]
For the right side:
$(x^{\prime}_{0}-x_{0})^{T}(\nabla Q(x_{0})-\nabla Q^{\prime}(x_{0}))$ $=(x^{\prime}_{0}-x_{0})^{T}(K-K^{\prime})x_{0}+(x^{\prime}_{0}-x_{0})^{T}(k^{\prime}-k)$
[…]
Then, the rest of the proof provided in [Daniel1973] still holds. Indeed, the propositions [Daniel1973, Propositions 3.1, 3.2] only tackle the constrained sets, and do not involve $k$ or $k^{\prime}$. Moreover, in, [Daniel1973, Section 4], the relationship between $k$ and $k^{\prime}$ is not used (except when using [Daniel1973, Theorem 2.1], which we generalized already); in fact, [Daniel1973, Section 4] only works on $Q(x)$. ∎
Proposition 3 (Convergence of SQP for upper-$C^{1}$ functions)
$\alpha x^{T}x\leq x^{T}H_{k}x\leq\beta x^{T}x$
(ii) For each $k$, there exists a KKT point of $Q(x_{k},H_{k},g_{x_{k}})$ (equation (10)) with a Lagrange multiplier vector $u_{k}$ such that $u_{k}||_{\infty}\leq r$.
$S^{0}(\bar{x})=\{p:c(\bar{x})+\nabla c(\bar{x})^{T}p<0\}\neq\emptyset$
is a KKT point of Problem (8).
Proof.
Once again, we follow the proof of [Han1997, Theorem 3.2] and adapt it to the case where f is not differentiable but only upper-$C^{1}$. For completeness, we reproduce the entire proof and indicate where our modified assumptions apply, while emphasizing that the proof itself is not our own.
By assumption (ii), $(p_{k},u_{k})$ is a KKT pair of $Q(x_{k},H_{k},g_{x_{k}})$ with $u_{k}||_{\infty}\leq r$. If $p_{k}=0$, then $(x_{k},u_{k})$ satisfies the KKT conditions (9) of Problem (8) and the algorithm can terminate at $x_{k}$.
Thus, for the following, we assume that $p_{k}\neq 0$. From Proposition 2.a) (originally [Han1997, Theorem 3.1], which we adapted to upper-$C^{1}$ functions), it is possible to choose $x_{k+1}$ as described in Algorithm 1, and in particular, it verifies:
$\theta_{r}(x_{k+1})<\theta_{r}(x_{k})+\epsilon_{k}\;.$
Let $\bar{x}$ be an accumulation point of $(x_{k})$ with $S^{0}(\bar{x})\neq\emptyset$. Without loss of generality (passing to the subsequence if necessary) we may assume
$x_{k}\rightarrow\bar{x}\;,H_{k}\rightarrow\bar{H}$
where the existence of $\bar{H}$ follows from assumption (i).
In the original proof [Han1997], the continuity of the gradient of $f$ ensures that $\nabla f(x_{k})$ converges to $\nabla f(\bar{x})$. Here, a similar property is maintained as follows. Since $f$ is locally Lipschitz, its subgradients $(g_{x_{k}})$ are uniformly bounded for $k$ sufficiently large, hence there exists a subsequence that converges to a vector $\bar{g}$. Without loss of generality (passing to the subsequence if necessary) we note
$g_{x_{k}}\rightarrow\bar{g}\;,$
and the upper semi-continuity of the Clarke subdifferential [Clarke1998, Proposition 1.5] yields $\bar{g}\in\partial f(\bar{x})$. Since $\bar{H}$ is positive definite and $S^{0}(\bar{x})\neq\emptyset$, the quadratic problem $Q(\bar{x},\bar{H},\bar{g})$ (originally $Q(\bar{x},\bar{H},\nabla f(\bar{x}))$ in [Han1997]) has a unique KKT point $\bar{p}$. If $\bar{p}=0$, then $\bar{x}$ is a KKT point of Problem (8) and the theorem follows. Thus, we now assume that $\bar{p}\neq 0$ and we will show that this leads to a contradiction.
From the continuity of $\nabla g(x)$, and by Lemmas 1 and 2, where $x_{k+1}$ and $x_{k}$ play the roles of $x_{0}$ and $x^{\prime}_{0}$, and $g_{x_{0}}$ and $g_{x^{\prime}_{0}}$ play the roles of $k$ and $k^{\prime}$, we deduce that:
$p_{k}\rightarrow\bar{p}$
In the original proof [Han1997], this follows [Daniel1973] (the equivalent to our Lemma 2) and the continuity of $\nabla f(x)$ instead (which is stronger than our Lemma 1).
Since $(u_{k})$ is uniformly bounded by $r$, it has an accumulation point $\bar{u}$. We note
$u_{k}\rightarrow\bar{u}$
with $\bar{u}||_{\infty}\leq r$. By passing to the limit in the KKT relation, which is valid for any $k$, we obtain that $\bar{u}$ is a Lagrange multiplier of $Q(\bar{x},\bar{H},\bar{g})$.
The rest of the proof of [Han1997] is unchanged, and is as follows. Let $\bar{\lambda}\in\begin{bmatrix}0&\lambda_{max}\end{bmatrix}$ be chosen such that
$\theta_{r}(\bar{x}+\bar{\lambda}\bar{p})=\min_{0\leq\lambda\leq\lambda_{max}}\;\theta_{r}(\bar{x}+\lambda\bar{p}).$
Let us set
$\gamma=\theta_{r}(\bar{x}+\bar{\lambda}\bar{p})-\theta_{r}(\bar{x})>0$
which is strictly positive by Proposition 2.a). Since $x_{k}+\bar{\lambda}p_{k}\rightarrow\bar{x}+\bar{\lambda}\bar{p}$, for sufficiently large $k$ we have:
$\theta_{r}(x_{k}+\bar{\lambda}p_{k})+\gamma/2<\theta_{r}(\bar{x})\;.$
However, for all $k$ we have $\theta_{r}(x_{k+1})<\theta_{r}(x_{k})+\epsilon_{k}$, and for sufficiently large $k$ we have
$\sum_{i=k}^{\infty}\epsilon_{i}<\gamma/2$
which leads to:
$\theta_{r}(\bar{x})$ $<\theta_{r}(x_{k+1}+\sum_{i=k+1}^{\infty}\epsilon_{i}$
$\leq\min_{0\leq\lambda\leq\lambda_{max}}\theta_{r}(x_{k}+\lambda p_{k})+\epsilon_{k}+\sum_{i=k+1}^{\infty}\epsilon_{i}$
$<\theta_{r}(x_{k}+\bar{\lambda}p_{k})+\gamma/2$
which contradicts the previous inequality $\theta_{r}(x_{k}+\bar{\lambda}p_{k})+\gamma/2<\theta_{r}(\bar{x})$. This proves that $\bar{p}=0$ and that $\bar{x}$ is a KKT point of Problem (8). ∎

### v4

@@ V-B Closed loop and control problem
The benchmark focuses on the attitude control of the satellite described above to follow a reference $r$. The control input $u$ of the plant $G(s,\delta)$ is the torque applied to the satellite (e.g. by reaction wheels, thrusters…) and the output vector contains the 3 angle measurements polluted by some noise $n$. The closed loop is represented in Figure 2, where $d(s)$ represents avionics: a 2-nd order Pade approximation of a 0.125s delay, and a 2-nd order low-pass filter with a cut-off frequency of 10 rad/s representing actuator dynamics. We note $\mathcal{K}$ the set of controllers $K(s)$ of the desired structure: one 4-th order controller per spacecraft axis, each one taking the measured angle and rate as inputs and generating a control torque, resulting in 24 decision parameters per controller (72 in total). The closed-loop plant has 50 states.
@@ V-C Worst-case search
In this section, we present the worst-case search for the spectral abscissa and the $\mathcal{H}_{\infty}$ norm, as presented in Section III. We selected three controllers $K_{1}$, $K_{2}$, $K_{3}$ obtained at early iterations of the robust controller tuning presented in Section V-D and analyze the robustness of the closed-loop system with three Test cases. For the sake of brevity, and since all 3 test cases yield similar conclusions, we present only the Test Case 1 in the body of the paper; Test Cases 2 and 3 are in Appendix B. Test case 1 illustrates how the proposed approaches perform in presence of several local optima, in different regions of the parametric space, that are not particularly rare. We search for $\max_{\delta\in\mathcal{D}}\;\alpha(A(K_{1},\delta))$.
The tests are ran on a 12th Gen Intel(R) Core(TM) i7-12700H 2.30 GHz, with 14 cores and 16GB of RAM. Computation times are provided for reference, but they may vary depending on hardware, software configuration, implementation details, optimization parameters, etc.
@@
The parameters of the Particle Swarm Optimization (PSO) were chosen as follows to favor global exploration over local exploitation. We set a range of 0.6 to 1.1 for the adaptive inertias (’InertiaRange’ option in MATLAB). Iterations terminate when performance is increased by less than 10% (”FunctionTolerance’) across 20 last iterations (’MaxStallIterations’). For a given computational budget, because PSO is sensitive to initialization [35], running several independent instances with different initializations can yield better results than increasing the number of particles or tightening the termination tolerance. Therefore, we performed 100 independent runs with the parameters given above, each with 500 particles (’SwarmSize’) – cf. also [43] for a discussion on population size in PSO. All other parameters were set to MATLAB’s default settings. Bearing in mind that our parameter choices were specifically tuned for multiple fast independent runs rather than to maximize the reliability of a single run, the results should be interpreted to determine how many such runs are required to reliably identify a worst case. We see that, with a few runs, we are able to consistently identify the worst-case configuration that was most often missed by the local SQP alone.
@@
As mentionned earlier, Appendix B presents two additional Test Cases with similar conclusions.
@@ V-D Robust control
In this section, we solve the robust control Problem (26) using Algorithm 2.
@@ VI Conclusion
To perform the worst-case search, we combined Particle Swarm Optimization (or Monte-Carlo sampling, if stochastic guarantees are desired) for global exploration and Sequential Quadratic Programming for local exploitation. Numerical experiments on a flexible spacecraft benchmark show that this method is able to reliably detect even rare worst-case configurations. Then, by iteratively constructing a set of active configurations, multi-objectives robust controller optimization was performed in reasonable runtime.
Finally, we emphasize that the proposed worst-case search, performing global exploration (albeit without deterministic guarantee of finding the global optimum) and being compatible with a large number of uncertain parameters and states (as a gradient-based method), has the potential to complement existing robust control techniques even in the unconstrained case.
@@ Acknowledgments
The authors thank Vincent Guibert for his feedback on the preprint. The authors acknowledge the use of an AI tool (ChatGPT: GPT-4) to assist in improving the wording and reformulation of selected sentences.